# Constraining the stochastic gravitational wave background using the future lunar seismometers

### v1

@@ Abstract
Motivated by the old idea of using the moon as a resonant gravitational-wave (GW) detector, as well as the recent updates in modeling the lunar response to GWs, we re-evaluate the feasibility of using a network of lunar seismometers to constrain the stochastic GW background (SGWB). In particular, using the updated model of the lunar response, we derive the pattern functions for the two polarizations of GW. With these pattern functions, we further calculate the overlap reduction functions for a network of lunar seismometers, where we have relaxed the conventional assumption that lunar seismometers are perfectly leveled to measure only the vertical acceleration. We apply our calculation to two future lunar projects, namely, Chang’e and the Lunar Gravitational-Wave Antenna (LGWA). We find that the two projects could constrain the SGWB to a level of $\Omega_{\text{GW}}^{\text{Chang'e}}<79$ and $\Omega_{\text{GW}}^{\text{LGWA}}<6.7\times 10^{-11}$, respectively. These results are better than the constraints placed previously on the SGWB in the mid-frequency band (around $10^{-3}-10~{}\text{Hz}$) by various types of experiments.
@@ I Introduction
Three recent progresses motivated us to revisit the calculation of the sensitivity of lunar seismometers to the SGWB. First, the physical interpretation and theoretical formulation of the lunar response to GW have evolved in recent years [30, 31, 32], and the response functions of the moon have been updated [31, 33, 34]. Conventionally, the lunar response to GW has been understood in two different ways, either from the tidal acceleration or from the local shear force induced by GW (the latter one corresponds to the result of an effective-field-theory approach [32]). In light of the recent studies, it becomes clear that the latter approach results in a response function which is directly proportional to the readout of a lunar seismometer. Second, deployment of new seismometers on the moon is imminent, including China’s Chang’e 7 and 8 projects [35, 36, 37] which are scheduled in 2026 and 2028, and NASA’s Farside Seismic Suite (FSS) scheduled in 2026 [38]. The India’s Chandrayan-3 project launched in 2023 also includes the Instrument for Lunar Seismic Activity (ILSA) [39]. Understanding their sensitivity to the SGWB is urgent. Third, depending on the design of seismometer and the deployment mechanism, future lunar seismometers may not be perfectly leveled, and one seismometer may be able to measure the accelerations in multiple directions, but the previous works commonly assumed that the seismometer arrays are responding only to the vertical acceleration. A related matter is that seismometers with different orientations are sensitive to different polarization states of the incoming GWs. But the previous works are mostly based on a single polarization state. Here, we aim to address the above issues.
@@ II.1 Surface response of the moon
the surface response of a radially heterogeneous elastic sphere can be written as:
@@
where we are evaluating the response at the location of $R\left(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta\right)$ ($R$ is the lunar radius). Following the definition in Yan24, $T_{r}$ and $T_{h}$ are the radial and horizontal response functions at the GW frequency $\omega_{g}/2\pi$. They are independent of the location of a detector or the propagation direction of GW. $\mathcal{Y}_{2m}$ is the real spherical harmonics, and $f^{m}$ is a function depending on the direction of the propagation of GW. The definitions of three base vectors in spherical coordinates are respectively
@@
Notice that the tensor $\mathbf{h}$ used above refers to a single polarization state with a single frequency. Therefore, the displacement solution, Eq. (2), although widely used in the past, applies only to a single polarization.
@@
$\mathbf{h}=\left(h_{+}\mathbf{e}_{+}+h_{\times}\mathbf{e}_{\times}\right)e^{i\omega_{g}t}~{},$ (5)
@@ II.2 Theory of detecting SGWB
where $\Delta f$ is the length of the integral interval. Therefore, the threshold $\Omega_{\text{GW}}$ that is possibly detectable would be
@@
By setting $\Delta f\simeq f$, the above equation can be further simplified to:
$\Omega_{\text{GW}}\simeq\frac{\eta}{H_{0}^{2}}\frac{f^{5/2}}{\sqrt{T}}\left(\frac{\sqrt{S_{n}}}{T_{r/h}}\right)^{2}$
$\simeq$ $3.45\times 10^{-8}\left(\frac{\eta}{10}\right)\left(\frac{H_{0}}{70\thinspace\text{km}\thinspace\text{s}^{-1}\thinspace\text{Mpc}^{-1}}\right)^{-2}\left(\frac{f}{1\thinspace\text{Hz}}\right)^{-3/2}$
@@
in which we have applied Eq. (7). The $f^{2}\sqrt{S_{n}}$ term approximately corresponds to the sensitivity of the seismometer, and $\eta$ is a factor of order unity coming from Eq. (12). Notice that this equation generally underestimates the detectability by one or two orders of magnitude, as we will discuss in Sec. IV.
@@ III.2 The ORF of an array of seismometers with arbitrary orientations
where $\zeta$ is a ratio defined by $T_{h}=\zeta T_{r}$. In this specific situation the maximum value of ORF depends on $r$. If $\zeta^{2}>1/3$, the maximum value is $58\zeta^{2}T_{r}^{2}/$ 5 when $B_{i}=C_{i}=1/\sqrt{2},~{}i=1,2$. Otherwise, if $\zeta^{2}<1/3$, the maximum becomes $158T_{r}^{2}/$ 15 when $B_{i}=C_{i}=0,~{}i=1,2$. For two seismometers $i$ and $j$ that are all horizontal (i.e., $A_{i}=A_{j}=0$), Eq. (24) can be further simplified to
@@
We note that these ORFs [Eq. (24) and Eq. (25)] may overestimate the SNR for the seismometer array in a small region, because two seismometers placed nearby could have gained correlated noise.
@@
$\Gamma^{\text{Chang'e}}\simeq\frac{8}{15}\sqrt{T_{r}^{4}+18T_{h}^{4}}~{},$ (28)
@@ IV Constraining the SGWB by lunar seismometers
Next, we calculate the constraint on the energy spectral density of the SGWB by Chang’e and LGWA. According to Eqs. (28) and (30) and the related discussions in Sec. III.2, we set $\Gamma^{\text{Chang'e}}=8\sqrt{T_{r}^{4}+18T_{h}^{4}}/15$ and $\Gamma^{\text{LGWA}}=16\sqrt{3}T_{h}^{2}/5$. Then the detection thresholds are derived in two different ways. (i) First, we use the approximation, Eq. (12), and set $\Delta f=f$, $\text{SNR}=1$, and $T=1\text{yr}$. The results are shown in Fig. 3 as the solid lines. (ii) Second, we perform a more accurate integration using Eq. (9), where we have assumed a flat spectrum for $\Omega_{\text{GW}}$ [i.e., setting $\Omega_{\text{GW}}(f)$ = $\Omega_{\text{GW}}$], and set $\text{SNR}=1$ and $T=1\text{yr}$. The integration is performed in the frequency domain of $10^{-3}-10~{}\text{Hz}$. The corresponding thresholds are shown in Fig. 3 as the horizontal dashed lines. The values corresponding to these two dashed lines are $\Omega_{\text{GW}}^{\text{Chang'e}}=79$ and $\Omega_{\text{GW}}^{\text{LGWA}}=6.7\times 10^{-11}$, respectively. It is worth mentioning that our approach can be easily extended to the cases with different shapes of SGWB energy spectrum, by changing the constant $\Omega_{\text{GW}}$ to a certain kind of function $\Omega_{\text{GW}}(f)$ in Eq. (9). In general, the sensitivity of LGWA might enable people to distinguish the differences between different SGWB models.
@@ V Summary and discussions
Using the pattern functions, we also constructed the ORFs for a network of two seismometers (Sec. III.1) as well as an array of an arbitrary number of seismometers (Sec. III.2). We applied our ORFs to two future projects, namely, Chang’e and LGWA (Sec. IV). We found that in the frequency band around $0.1$ Hz, the threshold SGWB detectable by Chang’e is two orders of magnitude better than the limit given previously by other missions. The sensitivity of LGWA would be even better, reaching a level as low as $\Omega_{\text{GW}}^{\text{LGWA}}\sim 6.7\times 10^{-11}$ throughout the mid-frequency band of $1010^{-3}-$ 10 Hz (Fig. 4).
Finally, we point out several caveats in this work which deserve future investigation. First, we have omitted the correlation between the noise of the seismometers, which will be crucial for the seismometers placed at close locations. This correlation may undermine the ability of a seismometer array in constraining the SGWB. Second, as we have mentioned in Yan24, we ignored the complex structure at the surface of the moon. This structure might influence the lunar response functions [34], and should be examined by the real data from future lunar seismometers. Third, a recent theoretical study of the interaction between GW and elastic body hints that the current model of the lunar response may be incomplete [32]. The difference between the results deserves further investigation.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203002, the Beijing Natural Science Foundation No. 1242018, and the National Natural Science Foundation of China (NSFC) Grants No. 11991053, No. 12073005, No. 12021003, No. 42325406. The authors would like to thank Jan Harms for useful discussions and comments.
@@
We first consider a new polarization state of GW that is different from Eq. (1). It can be written as

### v2

@@ Abstract
Motivated by the old idea of using the moon as a resonant gravitational-wave (GW) detector, as well as the recent updates in modeling the lunar response to GWs, we re-evaluate the feasibility of using a network of lunar seismometers to constrain the stochastic GW background (SGWB). In particular, using the updated model of the lunar response, we derive the pattern functions for the two polarizations of GW. With these pattern functions, we further calculate the overlap reduction functions for a network of lunar seismometers, where we have relaxed the conventional assumption that lunar seismometers are perfectly leveled to measure only the vertical acceleration. We apply our calculation to two future lunar projects, namely, Chang’e and the Lunar Gravitational-Wave Antenna (LGWA). We find that the two projects could constrain the SGWB to a level of $\Omega_{\text{GW}}^{\text{Chang'e}}<2.4\times 10^{2}$ and $\Omega_{\text{GW}}^{\text{LGWA}}<2.0\times 10^{-10}$, respectively, which corresponds to a signal-to-noise ratio of SNR $=3$. These results are better than the constraints placed previously on the SGWB in the mid-frequency band (around $10^{-3}-10~{}\text{Hz}$) by various types of experiments.
@@ I Introduction
Three recent progresses motivated us to revisit the calculation of the sensitivity of lunar seismometers to the SGWB. First, the physical interpretation and theoretical formulation of the lunar response to GW have evolved in recent years [30, 31, 32], and the response functions of the moon have been updated [31, 33, 34]. Conventionally, the lunar response to GW has been understood in two different ways, either from the tidal acceleration or from the local shear force induced by GW (the latter one corresponds to the result of an effective-field-theory approach [32]). In light of the recent studies, it becomes clear that the latter approach results in a response function which is directly proportional to the readout of a lunar seismometer. Second, deployment of new seismometers on the moon is imminent, including China’s Chang’e 7 and 8 projects [35, 36, 37] which are scheduled in 2026 and 2028, and NASA’s Farside Seismic Suite (FSS) scheduled in 2026 [38]. The India’s Chandrayan-3 project launched in 2023 also had the Instrument for Lunar Seismic Activity (ILSA) [39]. Understanding their sensitivity to the SGWB is urgent. Third, depending on the design of seismometer and the deployment mechanism, future lunar seismometers may not be perfectly leveled, and one seismometer may be able to measure the accelerations in multiple directions, but the previous works commonly assumed that the seismometer arrays are responding only to the vertical acceleration. A related matter is that seismometers with different orientations are sensitive to different polarization states of the incoming GWs. But the previous works are mostly based on a single polarization state. Here, we aim to address the above issues.
@@ II.1 Surface response of the moon
the surface response of a radially heterogeneous elastic sphere in the long-wavelength approximation can be written as:
@@
where we are evaluating the response at the location of $R\left(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta\right)$ ($R$ is the lunar radius). Following the definition in Yan24, $T_{r}$ and $T_{h}$ are the radial and horizontal response functions at the GW frequency $\omega_{g}/2\pi$. They are independent of the location of a detector or the propagation direction of GW. $\mathcal{Y}_{2m}$ is the real spherical harmonics, and $f^{m}$ is a function depending on the direction of the propagation of GW:
$f^{m}=4\sqrt{\frac{\pi}{15}}\times\left(\delta_{m,2}+\delta_{m,-2}\right)~{}.$
The definitions of three base vectors in spherical coordinates are respectively
@@
Notice that the tensor $\mathbf{h}$ used above refers to one single polarization state with a single frequency, which can also be represented by the commonly used polarization states $h_{+}$ and $h_{\times}$ as $h_{+}+h_{\times}$. Another polarization state $h_{+}-h_{\times}$ is discussed in the Appendix. Therefore, the displacement solution, Eq. (2), although widely used in the past, applies only to a single polarization.
@@
$\mathbf{h}$ $=\left(h_{+}\mathbf{e}_{+}+h_{\times}\mathbf{e}_{\times}\right)e^{i\omega_{g}t}$
$\mathbf{e}_{+}$ $=\begin{bmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{bmatrix}$
$\mathbf{e}_{\times}$ $=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{bmatrix}~{},$ (5)
@@ II.2 Theory of detecting SGWB
where $\Delta f$ is the bandwidth. Therefore, the threshold $\Omega_{\text{GW}}$ that is possibly detectable would be
@@
By setting $\Delta f\simeq f$ and $\Gamma\simeq\eta T_{r/h}^{2}$, the above equation can be further simplified to:
$\Omega_{\text{GW}}\simeq\frac{4\pi^{2}}{3\sqrt{2}}\frac{\text{SNR}}{\eta}\frac{1}{H_{0}^{2}}\frac{f^{5/2}}{\sqrt{T}}\left(\frac{\sqrt{S_{n}}}{T_{r/h}}\right)^{2}$
$\simeq$ $3.2\times 10^{-7}\left(\frac{\text{SNR}}{\eta}\right)\left(\frac{H_{0}}{70\thinspace\text{km}\thinspace\text{s}^{-1}\thinspace\text{Mpc}^{-1}}\right)^{-2}\left(\frac{f}{1\thinspace\text{Hz}}\right)^{-3/2}$
@@
in which we have applied Eq. (7). The $f^{2}\sqrt{S_{n}}$ term approximately corresponds to the sensitivity of the seismometer. Notice that this equation generally underestimates the detectability by one or two orders of magnitude, as we will discuss in Sec. IV.
@@ III.1 The ORF of two seismometers with special orientations
To better illustrate these configurations, we plot Fig. 1 to show the meanings of “ver”, “hor1” and “hor2”.
Figure 1: Schematic diagram showing different configurations of the readout directions of the seismometers on the moon. The grey arc represents the great circle crossing two seismometers.
@@ III.2 The ORF of an array of seismometers with arbitrary orientations
where $\zeta$ is a ratio defined by $T_{h}=\zeta T_{r}$. In this specific situation the maximum value of ORF depends on $\zeta$. If $\zeta^{2}>1/3$, the maximum value is $58\zeta^{2}T_{r}^{2}/$ 5 when $B_{i}=C_{i}=1/\sqrt{2},~{}i=1,2$. Otherwise, if $\zeta^{2}<1/3$, the maximum becomes $158T_{r}^{2}/$ 15 when $B_{i}=C_{i}=0,~{}i=1,2$. For two seismometers $i$ and $j$ that are all horizontal (i.e., $A_{i}=A_{j}=0$), Eq. (24) can be further simplified to
@@
We note that these ORFs [Eq. (24) and Eq. (25)] may overestimate the SNR for a seismometer array in a small region, because two seismometers placed nearby could have gained correlated environmental noise, e.g., from the seismic wave induced by a nearby source of moonquake. This noise will result in a worse sensitivity of the array to the GW background. However, this effect can be mitigated by adding another seismometer far from the previous two, e.g., on the far side of the moon as the phase 2 of the LGWA project proposes.
@@
$\Gamma^{\text{Chang'e}}$ $\simeq\sqrt{\left(\frac{8}{15}T_{r}^{2}\right)^{2}+\left(\frac{8}{5}T_{h}^{2}\right)^{2}+\left(\frac{8}{5}T_{h}^{2}\right)^{2}}$ (28)
$=\frac{8}{15}\sqrt{T_{r}^{4}+18T_{h}^{4}}~{},$
@@ IV Constraining the SGWB by lunar seismometers
Next, we calculate the constraint on the energy spectral density of the SGWB by Chang’e and LGWA. According to Eqs. (28) and (30) and the related discussions in Sec. III.2, we set $\Gamma^{\text{Chang'e}}=8\sqrt{T_{r}^{4}+18T_{h}^{4}}/15$ and $\Gamma^{\text{LGWA}}=16\sqrt{3}T_{h}^{2}/5$. Then the detection thresholds are derived in two different ways. (i) First, we use the approximation, Eq. (12), and set $\Delta f=f$, $\text{SNR}=3$, and $T=1\text{yr}$. The results are shown in Fig. 4 as the solid lines. (ii) Second, we perform a more accurate integration using Eq. (9), where we have assumed a flat spectrum for $\Omega_{\text{GW}}$ [i.e., setting $\Omega_{\text{GW}}(f)$ = $\Omega_{\text{GW}}$], and set $\text{SNR}=3$ and $T=1\text{yr}$. The integration is performed in the frequency domain of $10^{-3}-10~{}\text{Hz}$. The corresponding thresholds are shown in Fig. 4 as the horizontal dashed lines. The values corresponding to these two dashed lines are $\Omega_{\text{GW}}^{\text{Chang'e}}=2.4\times 10^{2}$ and $\Omega_{\text{GW}}^{\text{LGWA}}=2.0\times 10^{-10}$, respectively. It is worth mentioning that our approach can be easily extended to the cases with different shapes of SGWB energy spectrum, by changing the constant $\Omega_{\text{GW}}$ to a certain kind of function $\Omega_{\text{GW}}(f)$ in Eq. (9). In general, the sensitivity of LGWA might enable people to distinguish the differences between different SGWB models.
@@ V Summary and discussions
Using the pattern functions, we also constructed the ORFs for a network of two seismometers (Sec. III.1) as well as an array of an arbitrary number of seismometers (Sec. III.2). We applied our ORFs to two future projects, namely, Chang’e and LGWA (Sec. IV). We found that in the frequency band around $0.1$ Hz, the threshold SGWB detectable by Chang’e is two orders of magnitude better than the limit given previously by other missions. The sensitivity of LGWA would be even better, reaching a level as low as $\Omega_{\text{GW}}^{\text{LGWA}}\sim 2.0\times 10^{-10}$ throughout the mid-frequency band of $1010^{-3}-$ 10 Hz (Fig. 5).
Finally, we point out several caveats in this work which deserve future investigation. First, we have omitted the correlation between the noise of the seismometers, which will be crucial for the seismometers placed at close locations. This correlation may undermine the ability of a seismometer array in constraining the SGWB. Second, as we have mentioned in Yan24, we ignored the complex structure at the surface of the moon, which might lead to strong scattering of the seismic waves [45, 46]. This structure might influence the lunar response functions [34], and should be examined by the real data from future lunar seismometers. Third, a recent theoretical study of the interaction between GW and elastic body hints that the current model of the lunar response may be incomplete [32]. The difference between the results deserves further investigation.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China Grants No. 2021YFC2203002, No. 2023YFC2205800, the Beijing Natural Science Foundation Grant No. 1242018, the Fundamental Research Funds for the Central Universities Grant No. 310432103, and the National Natural Science Foundation of China (NSFC) Grants No. 11991053, No. 12073005, No. 12021003, No. 42325406. The authors would like to thank Jan Harms for helpful discussions.
@@
We first consider a new polarization state ($h_{+}-h_{\times}$) of GW that is different from Eq. (1). It can be written as